OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

SOOOO LOST

myininaya (myininaya):

what was the f(x) you got in 2?

OpenStudy (anonymous):

@myininaya I'll post what I have so far as a file.

myininaya (myininaya):

so you said f(x)=10x+3 so have you create h(x) yet?

myininaya (myininaya):

oh so you have h(x)=2x+10

OpenStudy (anonymous):

Yes and I'm stuck, I'll post the actual question also

myininaya (myininaya):

you found h(4) now find f(18) using f(x)=10x+6

OpenStudy (anonymous):

The year is 2033, and intelligent life has been found on Mars! Little green Martians have found their way to Earth and have appointed you as the Ultimate Math Ambassador to their planet. You must help them understand how Earthlings use math. Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make. Splott Fizzle 2x + 4 = –3x + 14 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 2x + 4 – 2x = –3x + 14 – 2x 5x + 4 = 14 4 = 14 – 5x 5x + 4 + 4 = 14 + 4 4 – 14 = 14 – 5x – 14 5x = 18 –12 = –5x 5 x divided by 5 is equal to 18 divided by 5 –12 + 5 = –5x + 5 x = 18 over 5 –7 = x Create your own function to teach the Martians about functions. Your function must contain at least two different operations. Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function. Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences. Using complete sentences, describe to the Martians how to find the inverse of your function. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

I'm just stuck on the last part

myininaya (myininaya):

pick a random value for x and you did which was 4 you evaluated h(4)=2(4)+10=18 but you need to evaluate f(h(4)) and h(f(4)) basically it looks like they want you to test the commutative property over operation composition You found h(4)=18 so what is f(h(4))=f(18)?

OpenStudy (anonymous):

So, if we substitute f(x) for f(18) it would work out like this: f(18)=10(18)+6= 180+6=186 ???

myininaya (myininaya):

right

myininaya (myininaya):

now you find h(f(4))

myininaya (myininaya):

and you want to see if f(h(x))=h(f(x)) for any value x.

OpenStudy (anonymous):

So which equation would I use? h(x)=2x+10 and substitute 18 for x?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

you are still using the same x value which is 4

myininaya (myininaya):

you are trying to see if f(h(4)) is the same as h(f(4))

myininaya (myininaya):

so you found f(h(4)) which was 186 now let's find h(f(4)) first find f(4)

OpenStudy (anonymous):

f(x)=10x+6 f(4)=10(4)+6 =46 I'm sorry, I'm horrible at math

myininaya (myininaya):

ok now we find h(46)

OpenStudy (anonymous):

h(x)=2x+10 =2(46)+10 =102

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

right so clearly f(h(x)) does not equal h(f(x)) for all x. What would we need in order for f(g(x)) to equal g(f(x)) for all x over the domain of f composed of g or g composed of f? what do f and g have to be?

OpenStudy (anonymous):

I literally do not have a clue, how did g come in? Lol I feel so dumb

myininaya (myininaya):

I was just introducing a new function and it was also a hint

myininaya (myininaya):

because you used g somewhere on your paper

OpenStudy (anonymous):

Right, as the inverse

myininaya (myininaya):

so if f and g are inverses then f(g(x)) will always equal g(f(x)) for any value x right? but h was not the inverse of f so f(h(x)) will not always equal h(f(x)) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

yes?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So I'm done then, right? Because the last question was if they would always equal and the answer to that is no.

myininaya (myininaya):

right they would equal for all x in the domain if f and h were inverses

myininaya (myininaya):

but they aren't

myininaya (myininaya):

or if the two functions were the same function :)

myininaya (myininaya):

like f(f(x))=f(f(x)) for all x right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, is all of this correct?

myininaya (myininaya):

yeah looks good... I would just change those 3 x's down there to 46 to show you are finding the value of h(46) and i know you are because you said you are but it just looks better to say h(46)=2(46)+10 h(46)=92+10 h(46)=102 well and also close to the start of number 6 I would say h(4)=2(4)+10=8+10=18 those are the only changes to number 6 I would make.

myininaya (myininaya):

and I guess you conclude somewhere after finding h(46)=102 that h(f(4))=102

OpenStudy (anonymous):

Thank you soooo much! I would hug you if I could! I hate math lol

myininaya (myininaya):

You did good.

OpenStudy (anonymous):

You're awesome!